We are having a hard time solving this equation 17=4(a-2)+2a
Answer by Jk22(389)

17=4(a-2)+2a=4a-8+2a (distributivity)
17=6a-8 (collect same monomial)
17+8=6a (+8, regularity)
25=6a (divide through 6)
a=25/6
Verification : a-2=(25-12)/6=13/6, 2*a=50/6, 4(a-2)+2a=(4*13+50)/6=102/6
6*17=60+6*7=60+42=102, hence 102/6=17.